# Robustness of Decentralised Learning to Nodes and Data Disruption

### v1

@@ Abstract
In the vibrant landscape of AI research, decentralised learning is gaining momentum. Decentralised learning allows individual nodes to keep data locally where they are generated and to share knowledge extracted from local data among themselves through an interactive process of collaborative refinement. This paradigm supports scenarios where data cannot leave local nodes due to privacy or sovereignty reasons or real-time constraints imposing proximity of models to locations where inference has to be carried out. The distributed nature of decentralised learning implies significant new research challenges with respect to centralised learning. Among them, in this paper, we focus on robustness issues. Specifically, we study the effect of nodes’ disruption on the collective learning process. Assuming a given percentage of “central” nodes disappear from the network, we focus on different cases, characterised by (i) different distributions of data across nodes and (ii) different times when disruption occurs with respect to the start of the collaborative learning task. Through these configurations, we are able to show the non-trivial interplay between the properties of the network connecting nodes, the persistence of knowledge acquired collectively before disruption or lack thereof, and the effect of data availability pre- and post-disruption. Our results show that decentralised learning processes are remarkably robust to network disruption. As long as even minimum amounts of data remain available somewhere in the network, the learning process is able to recover from disruptions and achieve significant classification accuracy. This clearly varies depending on the remaining connectivity after disruption, but we show that even nodes that remain completely isolated can retain significant knowledge acquired before the disruption.
@@ 1 Introduction
Under these assumptions, we analyse the robustness of decentralised learning under three configurations, which we denote as Case 1, Case 2, and Case 3, respectively (see Section 4 for more details), characterised by increased levels of disruptions. In Case 1, after an initial period where the decentralised process proceeds without disruptions, we assume that a certain percentage of “central” nodes disappear. However, these nodes do not hold any local data, so the effect is a disruption of the network structure only. In Case 2, we assume that the same nodes also hold local data, and therefore, the disruption consists of the loss of both data and connectivity. Finally, in Case 3, we assume that central nodes hold a significantly higher share of data with respect to the other nodes. Therefore, Case 3 represents the loss of connectivity and a particularly extended set of data on which pre-disruption models could have been trained. In all these cases, we analyse the performance in terms of accuracy achieved by surviving nodes after a certain number of training rounds. As the considered disruptions partition the network into smaller disconnected subgraphs, we also analyse the performance inside the different clusters of connected nodes that are generated after disruption, starting from isolated nodes to the largest connected component. In all cases, we compare the accuracy achieved by the surviving nodes after the disruption in comparison to the baseline case where no disruption occurs and the case when disruption occurs at time 0. The latter represents the boundary case where surviving nodes cannot exploit any collaborative learning process before the disruption.
@@ 2 Related work
As stated in the introduction, network disruption is a central concern in today’s interconnected world, whether the network is a communication network, a transportation network, or a social network. On the other hand, complex networks serve as the backbone of numerous real-world systems, from biological processes and social interactions to technological infrastructures like the internet and power grids[4]. As a result, the study of complex network models undergoing network disruption has attracted a lot of attention in the field of network science. Understanding the vulnerability and resilience of complex networks is essential for developing systems that can withstand failures and disruptions. This is particularly important in disaster management, where the goal is to minimise the impact of critical events on infrastructures and services. Different network topologies exhibit varying levels of resilience and vulnerability. For example, a centralised network is highly vulnerable to single points of failure, since all network communication flows through a central node or hub. Therefore, this topology is particularly vulnerable to disruptions that target the central node.
Within the context of complex networks and their applications to real-world scenarios, networks’ structures such as the Barabasi-Albert model, are frequently used as reference models. This is because of their ability to capture the inhomogeneous connectivity distribution characteristic of many real-world networks, thanks to the scale-free property exhibited by this model. In [5], the authors investigated the tolerance of complex networks with scale-free properties to random errors and targeted attacks. They looked at how the diameter of the networks changed when a small number of nodes were removed, disrupting the interconnectedness of the system, in order to address the networks’ fault tolerance. They found that scale-free networks display high resilience against random failures, largely due to their heterogeneous topology. This resilience stems from the vast majority of nodes having few connections, making it statistically unlikely for a random failure to hit a hub, which is crucial for maintaining the network’s connectivity. On the other hand, in a related study [6], the authors examined the vulnerability of different network models, including Barabasi-Albert graphs, to intentional attacks. Their investigation involved an analysis of how scale-free networks respond to various attack strategies that specifically target nodes based on their properties, such as degree and centrality measures. By examining the behaviour and robustness of these networks under different attack scenarios, they provided valuable insights into the dynamics of scale-free networks and their susceptibility to intentional attacks. They found that these network structures are vulnerable to targeted attacks on their hubs.
@@
Decentralised Learning (DL) extends the typical settings of Federated Learning [3] by removing the existence of the central parameter server. This approach is gaining momentum as it merges the privacy benefits inherent in Federated Learning with the capabilities of decentralised and uncoordinated optimisation and learning processes. In [10], a DL model is deployed for a healthcare application, enabling multiple hospitals to collaboratively train a Neural Network model while keeping their data private. Another work [11] introduces a federated consensus algorithm that extends the FedAvg method proposed by [3] for use in decentralised environments, with emphasis on industrial and IoT applications. The work in [12] proposes a Federated Decentralized Average based on SGD in their research, where they incorporate a momentum term to counterbalance potential drift caused by multiple updates, along with a quantisation strategy to reduce communication requirements. In [13], the authors introduce a novel method for collaborative training in DL environments by means of a secret sharing schema and a multi-party average computation, aiming to preserve privacy and security. The authors of [14] explore the influence of the underlying network architecture on the learning efficacy in a fully decentralised learning system, while [15] investigates the impact of network topology on the optimal model initialisation in decentralised settings. However, the issues surrounding system robustness remain unaddressed within the domain of DL. However, the effectiveness and reliability of DL systems are crucial for their success. Our research represents a step towards filling this gap.
@@
This paper extends the work in [16] along several directions. Firstly, we employ a new strategy for simulating disruptions, ensuring fairer comparisons among different learning configurations by imposing disruptions once the average accuracy of the system reaches a specific target threshold. This contrasts with the previous approach, where disruptions occurred at fixed intervals. Furthermore, a novel scenario (Case 3) has been introduced, detailing situations where disrupted nodes exhibit a disproportionate proficiency in classifying specific classes of images. This setup enables evaluating surviving nodes’ ability to retain valuable knowledge provided by disrupted nodes after disconnection. Moreover, the paper now includes an analysis of the impact of different centrality measures for selecting nodes to be switched off, as well as a percolation analysis to characterise changes in network connectivity as nodes are removed, detailed in Section 4.2. Additionally, for Case 1 and Case, we now consider a more challenging (from the learning standpoint) data distribution. Finally, a comparison of the performance of isolated nodes versus the largest cluster has been added to illustrate the robustness of the decentralised learning process in extreme cases following disruption.
@@ 3.1 Decentralized learning
We represent the network connecting the nodes as a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}$ is the set of edges. The decentralised learning algorithms designed to operate on a network of nodes are typically composed of two main blocks: one for the local training of the model using local data and the other one devoted to the exchange and aggregation of the models’ updates. These operations are executed by each node, atomically, within a single communication round. Each node $i\in\mathcal{V}$ has a local training dataset $\mathcal{D}_{i}$ (containing tuples of features and labels $(\mathbf{x},y)\in\mathcal{X}\times\mathcal{Y}$) and a local model $h$ defined by weights $\mathbf{w}_{i}$, such that $h(\mathbf{x};\mathbf{w}_{i})$ yields the prediction of label $y$ for input $\mathbf{x}$. Let us denote with $\mathcal{D}=\bigcup_{i}\mathcal{D}_{i}$ and with $\mathcal{P}$ the label distribution in $\mathcal{D}$. In this paper, we consider both the case where $\mathcal{P}_{i}$ (i.e., the label distribution of the local dataset on node $i$) is different from $\mathcal{P}$, as well as the case where label distributions $\mathcal{P}_{i}$ are IID (independent and identically distributed) across all devices (hence, $\mathcal{P}_{i}\sim\mathcal{P}$). We also assume homogeneous initialisation of local models across devices, i.e., we assume that, at time 0, $\mathbf{w}_{i}=\mathbf{w}_{j}$ for any pair of nodes $i,j$ holds true. The effect of heterogenous initialisation on decentralised learning is investigated in [17], while in this paper we focus on the effect of disruption on the learning process, which is an orthogonal problem.
@@ 3.2 Network resilience and disruption strategy
This is how we use this methodology in this paper. Specifically, as explained in Section 4.2, we carry out a preliminary percolation analysis to identify the percentage of nodes to be disrupted to generate an interesting mix of connected components, from large ones to isolated ones. Then, we study the evolution of decentralised learning when such a percentage of nodes is disrupted by varying the point in time when these disruptions occur (to give more or less time to the underlying decentralised learning process to work on the original network before disruption). Specifically, we characterise the evolution of the learning process in the different components that emerge after disruption to show the interplay between the quality of the model learned before disruption, the residual connectivity, and the presence of data (or lack thereof) in the surviving components.
@@ 4.1 Communication network topology
The topology of the Barabasi-Albert graph is considered able to capture important features of real-life networks by incorporating two fundamental principles: preferential attachment (new nodes in the network tend to connect to existing nodes that already have a high number of connections) and organic growth (which aligns with the dynamic nature of many real-world systems). By reproducing these principles, a BA graph model successfully emulates the power-law degree distribution observed in numerous networks, such as social networks, the World Wide Web, and collaboration networks. For this reason, it makes sense to consider it a realistic topology for our network graph.
Figure 1: Barabasi-Albert Network before (left) and after (right) the disruption. The colouring of nodes is based on the size of the cluster they belong to after disruption. The black nodes in the left image are the nodes that will be switched off when the disruption condition is fulfilled.
4.2 Selection of “disrupted” nodes
@@
In Figure 2(a), we show the behaviour of the ratio $\Phi$ in our network for different centrality measures: structural hole score [19], betweenness centrality and degree centrality. Structural holes refer to the gaps or “holes” in a network where there are no direct connections between two or more groups of people. These gaps represent opportunities for nodes to act as intermediaries or brokers by connecting otherwise disconnected groups and facilitating the flow of information, resources, or ideas between them. Thus, the structural hole score quantifies the importance of a node in promoting information flow within the network. Nodes with higher structural hole scores are crucial in facilitating communication across different groups. The betweenness centrality quantifies the extent to which a node lies on the shortest paths between pairs of other nodes in the network. Hence, it measures the importance of a node in connecting other nodes. The degree centrality measures the importance of a node within a network based on the number of connections it has to other nodes. We refer the interested reader to [2] for a formal definition of these centrality metrics.
@@
We progressively remove those nodes with the highest value of the selected centrality measure. This is effectively equivalent to analysing the robustness of the Barabasi-Albert graph undergoing targeted attacks towards the most central nodes. As we can see in Figure 2(a), for the first 5% of nodes removed (exactly 5 nodes in our case), the impact of all centrality measures is equal. After removing the initial 5% of nodes, the degree centrality and structural hole score behave similarly, whereas the betweenness centrality curve remains generally higher. This implies that a larger portion of the network survives w.r.t. the other centralities, meaning that the graph is more robust to node removal when a targeted attack considers betweenness centrality. The reason why the curves behave the same for the first 5 nodes is that their centrality values rank the same nodes regardless of the specific metric considered, as it can be seen in Figure 2(c). In general, the figure shows that all the centrality measures have a linear directly proportional relationship. This means that nodes with high degrees also exhibit high betweenness centrality and high structural hole score. While the relationship remains directly proportional, subtle variations appear beyond the top five nodes. Other important factors in a disruption analysis are the number and size of the connected components that survive after the disruption. As we can see in Figure 2(b), after the disruption, there will be many isolated nodes (component size 1) and few connected components: the larger the size, the smaller the number. Note that the effect of disruption under degree centrality and structural hole score results in a higher number of isolated nodes than under the betweenness centrality case, meaning that undergoing such disruption is more detrimental to the network as it isolates more nodes.
@@
Figure 2(a) shows that the biggest drop in the largest connected component size happens at around 10% node removal, with a similar pattern for all centrality measures. Therefore, we decided to remove the 10% most central nodes. Due to its importance in measuring the importance of a node in promoting information flow, we select the structural hole score as our centrality metric. Note, though, that we are effectively also removing exactly the 10% of nodes with the highest degree centrality, as can be seen in Figure 2(a).
@@
4.3 Time of disruption
In disaster scenarios, disruptions present themselves in unpredictable ways. They can happen as soon as the collaborative learning process starts, in the middle, or when it is approaching its end. The loss of nodes is expected to have a different impact depending on the current progress of the learning process: weak when it happens in the final phase, stronger early on. In the latter case, the key question is whether the decentralised learning process is able to recover and catch up over time. As we discuss later on in Section 5.1, different learning configurations might exhibit different temporal dynamics, with one being slower than another one, even if ultimately their accuracy converges to the same value. To account for these different learning speeds and enable fair comparisons among the different scenarios considered, we thus force disruptions to happen once the average accuracy of the system has reached a certain target threshold. For example, in Section 5.2, we consider disruptions happening when the average accuracy in the system has reached 0.7, 0.75, 0.8. This guarantees that disruptions affect systems that are similarly “skilled” in the learning process at the time of disruption.

### v2

@@ Abstract
In the active landscape of AI research, decentralised learning is gaining momentum. Decentralised learning allows individual nodes to keep data locally where they are generated and to share knowledge extracted from local data among themselves through an interactive process of collaborative refinement. This paradigm supports scenarios where data cannot leave the data owner node due to privacy or sovereignty reasons or real-time constraints imposing proximity of models to locations where inference has to be carried out. The distributed nature of decentralised learning implies significant new research challenges with respect to centralised learning. Among them, in this paper, we focus on robustness issues. Specifically, we study the effect of nodes’ disruption on the collective learning process. Assuming a given percentage of “central” nodes disappear from the network, we focus on different cases, characterised by (i) different distributions of data across nodes and (ii) different times when disruption occurs with respect to the start of the collaborative learning task. Through these configurations, we are able to show the non-trivial interplay between the properties of the network connecting nodes, the persistence of knowledge acquired collectively before disruption or lack thereof, and the effect of data availability pre- and post-disruption. Our results show that decentralised learning processes are remarkably robust to network disruption. As long as even minimum amounts of data remain available somewhere in the network, the learning process is able to recover from disruptions and achieve significant classification accuracy. This clearly varies depending on the remaining connectivity after disruption, but we show that even nodes that remain completely isolated can retain significant knowledge acquired before the disruption.
@@ 1 Introduction
Under these assumptions, we analyse the robustness of decentralised learning under three configurations, which we denote as Case 1, Case 2, and Case 3, respectively (see Section 4 for more details), characterised by increased levels of disruptions. In Case 1, after an initial period where the decentralised process proceeds without disruptions, we assume that a certain percentage of “central” nodes disappear. However, these nodes do not hold any local data, so the effect is a disruption of the network structure only. In Case 2, we assume that the same nodes also hold local data, and therefore, the disruption consists of the loss of both data and connectivity. Finally, in Case 3, we assume that central nodes hold a significantly higher share of data with respect to the other nodes. Therefore, Case 3 represents the loss of connectivity and a particularly extended set of data on which pre-disruption models could have been trained. In all these cases, we analyse the performance in terms of accuracy achieved by surviving nodes after a certain number of training rounds. As the considered disruptions partition the network into smaller disconnected subgraphs, we also analyse the performance inside the different connected components of connected nodes that are generated after disruption, starting from isolated nodes to the largest connected component. In all cases, we compare the accuracy achieved by the surviving nodes after the disruption in comparison to the baseline case where no disruption occurs and the case when disruption occurs at time 0. The latter represents the boundary case where surviving nodes cannot exploit any collaborative learning process before the disruption.
@@ 2 Related work
2.1 Decentralised learning
Decentralised Learning (DL) extends the typical settings of Federated Learning (FL) [3] by removing the existence of the central parameter server. This approach is gaining momentum as it merges the privacy benefits inherent in Federated Learning with the capabilities of decentralised and uncoordinated optimisation and learning processes. In [4], a DL model is deployed for a healthcare application, enabling multiple hospitals to collaboratively train a Neural Network model while keeping their data private. Another work [5] introduces a federated consensus algorithm that extends the FedAvg method proposed by [3] for use in decentralised environments, with emphasis on industrial and IoT applications. The work in [6] proposes a Federated Decentralized Average based on SGD in their research, where they incorporate a momentum term to counterbalance potential drift caused by multiple updates, along with a quantisation strategy to reduce communication requirements. In [7], the authors introduce a novel method for collaborative training in DL environments by means of a secret sharing schema and a multi-party average computation, aiming to preserve privacy and security. The authors of [8] explore the influence of the underlying network architecture on the learning efficacy in a fully decentralised learning system. As discussed later in the paper, in this work we focus on DecAvg, a learning strategy that extends the widely-used FedAvg [3] to the decentralized domain, and generalises similar strategies proposed in [5, 6]. We deliberately chose not to explore more complex policies, as our primary objective is to examine the effects of network and data disruptions. Introducing more sophisticated approaches could have obscured the fundamental insights into these phenomena.
In addition to the algorithmic contributions discussed above, there have been efforts to analytically characterize decentralized learning in relation to graph topology. In [9], the authors present a comprehensive analytical framework that unifies various decentralized SGD methods. This framework supports local SGD updates, synchronous updates, and pairwise gossip updates within a network topology, and offers universal convergence rates for solving both convex and non-convex problems. Notably, their analysis operates under less stringent assumptions about network properties compared to earlier research. Similarly, in [10], authors examine the generalization and stability of decentralized SGD, providing a theoretical framework that correlates the network’s topology with the generalization of decentralized SGD. Their main focus is on how the spectral gap affects learning outcomes. Their analysis assumes IID data distribution and simple network topologies such as rings, grids and complete networks. Finally, [11] investigates the impact of network topology on the optimal model initialisation in decentralised settings, proposing a novel initialization strategy that speeds up the learning convergence in the initial phase.
2.2 Robustness of complex networks
@@
As stated in the introduction, network disruption is a central concern in today’s interconnected world, whether the network is a communication network, a transportation network, or a social network. On the other hand, complex networks serve as the backbone of numerous real-world systems, from biological processes and social interactions to technological infrastructures like the internet and power grids [12]. As a result, the study of complex network models undergoing network disruption has attracted a lot of attention in the field of network science. For a recent review of the issues of robustness in complex networks, we refer the interested reader to [13]. Understanding the vulnerability and resilience of complex networks is essential for developing systems that can withstand failures and disruptions. This is particularly important in disaster management, where the goal is to minimise the impact of critical events on infrastructures and services. Different network topologies exhibit varying levels of resilience and vulnerability. For example, a centralised network is highly vulnerable to single points of failure, since all network communication flows through a central node or hub. Therefore, this topology is particularly vulnerable to disruptions that target the central node.
Within the context of complex networks and their applications to real-world scenarios, networks’ structures such as the Barabasi-Albert model, are frequently used as reference models. This is because of their ability to capture the inhomogeneous connectivity distribution characteristic of many real-world networks, thanks to the scale-free property exhibited by this model. In [14], the authors explored how scale-free networks respond to random failures by examining changes in network diameter when a few nodes are removed. Their findings revealed that these networks are remarkably resilient to random failures, as the chances of a random event affecting one of the critical hubs is low due to the network’s heterogeneous topology. However, this resilience does not extend to deliberate attacks targeting the network hubs (high degree nodes). Random networks (such as the Erdös-Rényi graph), on the other end, are typically very robust to targeted attacks but vulnerable to random attacks, because, due their connections being more evenly distributed, the loss of any node can affect network connectivity more uniformly. In [15], the authors analyzed the vulnerability of scale-free networks, such as Barabasi-Albert graphs, to targeted attacks, where nodes are removed based on properties like degree or centrality. Their results showed that these networks are highly vulnerable when hubs are specifically targeted, as the removal of just a few key nodes can severely disrupt the network’s overall connectivity. This highlights a significant trade-off in scale-free networks: while they handle random failures well, they are critically susceptible to intentional attacks on their most connected nodes.
2.3 Robustness of federated learning
@@
In the literature, the term “robustness” is sometimes used to describe both resilience to failures and security threats posed by malicious actors. However, in this work, we distinguish between the two: we use “robustness” specifically to refer to resilience against failures. The security of federated learning has received significant attention in recent years. For an in-depth overview of the current state of security and privacy challenges in FL, we refer the reader to [19]. The authors of this survey highlight key security concerns, particularly targeted attacks from malicious actors, such as data poisoning and model poisoning, which can undermine the integrity of FL systems.
2.4 Robustness of decentralised learning
While federated learning imposes a star topology, where all nodes communicate through a central server, decentralized learning removes such constraints, allowing any network topology among the nodes. As a result, decentralized learning is shaped by the structure and dynamics of the underlying complex system that governs node interactions, as well as the specifics of the learning process itself. As DL is a relatively novel research area, both its security and its robustness remain underexplored in the literature. Regarding the topic of security in DL, authors in [20] explore the models’ robustness in decentralized learning against poisoning attacks, which can degrade model performance. It introduces DART, a tool for analysing DL model resilience to security threats, capable of simulating various attacks and evaluating defence mechanisms. The study compares centralized and decentralized paradigms under different attack scenarios and examines the effectiveness of defences. Authors in [20] conclude that the research landscape addressing security issues in FL/DL is primarily focused on FL, with fewer studies addressing purely DL, leaving the issues surrounding system security in DL remain largely unaddressed.
@@ 2.4 Robustness of decentralised learning
From the robustness standpoint, which is the focus of this work, the closest contributions are  [21] and [22]. The authors in [21] took into account decentralized FL where machine learning models are trained on edge devices using unreliable communication protocols, specifically UDP. They employed a modified version of traditional DSGD, which they call Soft-DSGD that updates model parameters with partially received data. The authors show that Soft-DSGD achieves the same convergence rate as standard decentralized methods while being more communication-efficient. Numerical results highlight its improved performance in scenarios with unreliable networks. However, in their setting, they account for partial communication loss and do not address the issue of client failure. To the best of our knowledge, the only work that focuses on node failures in DL from a topological perspective is [22]. This paper extends our work in [22] in several directions. Firstly, we employ a new strategy for simulating disruptions, ensuring fairer comparisons among different learning configurations by imposing disruptions once the average accuracy of the system reaches a specific target threshold. This contrasts with the previous approach, where disruptions occurred at fixed intervals. Furthermore, a novel scenario (Case 3) has been introduced, detailing situations where disrupted nodes exhibit a disproportionate proficiency in classifying specific classes of images. With this setup, we evaluate the ability of surviving nodes to retain valuable knowledge provided by disrupted nodes after disconnection. Moreover, the paper now includes an analysis of the impact of different centrality measures for selecting nodes to be switched off, as well as a percolation analysis to characterise changes in network connectivity as nodes are removed, detailed in Section 4.2. Additionally, for Case 1 and Case 2, we now consider a more challenging (from the learning standpoint) data distribution. Finally, a comparison of the performance of isolated nodes versus the largest connected component has been added to illustrate the robustness of the decentralised learning process in extreme cases following disruption.
@@ 3.1 Decentralized learning
We represent the network connecting the nodes as a graph $\mathcal{G}(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ denotes the set of nodes and $\mathcal{E}$ is the set of edges. The decentralised learning algorithms designed to operate on a network of nodes are typically composed of two main blocks: one for the local training of the model using local data and the other one devoted to the exchange and aggregation of the models’ updates. These operations are executed by each node, atomically, within a single communication round. Each node $i\in\mathcal{V}$ has a local training dataset $\mathcal{D}_{i}$ (containing tuples of features and labels $(\mathbf{x},y)\in\mathcal{X}\times\mathcal{Y}$) and a local model $h$ defined by weights $\mathbf{w}_{i}$, such that $h(\mathbf{x};\mathbf{w}_{i})$ yields the prediction of label $y$ for input $\mathbf{x}$. Let us denote with $\mathcal{D}=\bigcup_{i}\mathcal{D}_{i}$ and with $\mathcal{P}$ the label distribution in $\mathcal{D}$. In this paper, we consider both the case where $\mathcal{P}_{i}$ (i.e., the label distribution of the local dataset on node $i$) is different from $\mathcal{P}$, as well as the case where label distributions $\mathcal{P}_{i}$ are IID (independent and identically distributed) across all devices (hence, $\mathcal{P}_{i}\sim\mathcal{P}$). We also assume homogeneous initialisation of local models across devices, i.e., at time 0, $\mathbf{w}_{i}=\mathbf{w}_{j}$ for any pair of nodes $i,j$. The effect of heterogenous initialisation on decentralised learning is investigated in [23], while in this paper we focus on the effect of disruption on the learning process, which is an orthogonal problem.
@@ 3.2 Network resilience and disruption strategy
This is how we use this methodology in this paper. Specifically, as explained in Section 4.2.1, we carry out a preliminary percolation analysis to identify the percentage of nodes to be disrupted to generate an interesting mix of connected components, from large ones to isolated ones. Then, we study the evolution of decentralised learning when such a percentage of nodes is disrupted by varying the point in time when these disruptions occur (to give more or less time to the underlying decentralised learning process to work on the original network before disruption). Specifically, we characterise the evolution of the learning process in the different components that emerge after disruption to show the interplay between the quality of the model learned before disruption, the residual connectivity, and the presence of data (or lack thereof) in the surviving components.
@@ 4.1 Communication network topology
The topology of the Barabasi-Albert graph is considered able to capture important features of real-life networks by incorporating two fundamental principles: preferential attachment (new nodes in the network tend to connect to existing nodes that already have a high number of connections) and organic growth (which aligns with the dynamic nature of many real-world systems). By reproducing these principles, a BA graph model successfully emulates the power-law degree distribution observed in numerous networks, such as social networks, the World Wide Web, and collaboration networks. For this reason, it makes sense to consider it a realistic synthetic topology for our network graph.
The analysis carried out in this paper has also been performed on the much-studied “Zachary’s Karate Club” graph [25]. It is a real-life network of 34 nodes corresponding to members of a sports club with links representing social relationships. The network is known to have a two-community structure documented by the separation of the club into two parts after a conflict between two instructors in the club. We selected this graph because it is based on real data and because it features a community structure that is absent in BA topologies. The results obtained for the Karate Club graph well align to the ones obtained for the BA network: it seems that the removal of highly central nodes destroys the community structure of the Karate Club graph, so that it does not have an impact on the learning process. For this reason, we decided to leave the discussion of the results related to this topology to B and keep the paper focused on BA.
Figure 1: Barabasi-Albert Network before (left) and after (right) the disruption. The colouring of nodes is based on the size of the connected component they belong to after disruption. The black nodes in the left image are the nodes that will be switched off when disruption takes place.
4.2 Modelling disruption
In this section, we discuss how we select (i) the nodes to be disconnected from the network and (ii) the time when disruption occurs.
4.2.1 Selection of “disrupted” nodes
@@
In Figure 2(a), we show the behaviour of the ratio $\Phi$ in the BA network for different centrality measures: structural hole score [26], betweenness centrality and degree centrality. Structural holes refer to the gaps or “holes” in a network where there are no direct connections between two or more groups of people. These gaps represent opportunities for nodes to act as intermediaries or brokers by connecting otherwise disconnected groups and facilitating the flow of information, resources, or ideas between them. Thus, the structural hole score quantifies the importance of a node in promoting information flow within the network. Nodes with higher structural hole scores are crucial in facilitating communication across different groups. The betweenness centrality quantifies the extent to which a node lies on the shortest paths between pairs of other nodes in the network. Hence, it measures the importance of a node in connecting other nodes. The degree centrality measures the importance of a node within a network based on the number of connections it has to other nodes. We refer the interested reader to [2] for a formal definition of these centrality metrics.
@@
We progressively remove those nodes with the highest value of the selected centrality measure. This is effectively equivalent to analysing the robustness of the Barabasi-Albert graph undergoing targeted attacks towards the most central nodes. As we can see in Figure 2(a), for the first 5% of nodes removed (exactly 5 nodes in our case), the impact of all centrality measures is equal. After removing the initial 5% of nodes, the degree centrality and structural hole score behave similarly, whereas the betweenness centrality curve remains generally higher. This implies that a larger portion of the network survives w.r.t. the other centralities, meaning that the graph is more robust to node removal when a targeted attack considers betweenness centrality. The reason why the curves behave the same for the first 5 nodes is that their centrality values rank the same nodes regardless of the specific metric considered, as it can be seen in Figure 2(c). In general, the figure shows that all the centrality measures have a linear directly proportional relationship. This means that nodes with high degrees also exhibit high betweenness centrality and high structural hole score. While the relationship remains directly proportional, subtle variations appear beyond the top five nodes. Other important factors in a disruption analysis are the number and size of the connected components that survive after the disruption. As we can see in Figure 2(b), after the disruption, there will be many isolated nodes (a component of size 1) and few connected components: the larger the size, the smaller the number. Note that the effect of disruption under degree centrality and structural hole score results in a higher number of isolated nodes than under the betweenness centrality case, meaning that undergoing such disruption is more detrimental to the network as it isolates more nodes.
@@
Figure 2(a) shows that the biggest drop in the largest connected component size happens at around 10% node removal, with a similar pattern for all centrality measures. Therefore, we decided to remove the 10% most central nodes. Due to its importance in measuring the importance of a node in promoting information flow, we select the structural hole score as our centrality metric. Note, though, that we are effectively also removing exactly the 10% of nodes with the highest degree centrality, as seen in Figure 2(a).
In this work, we specifically address the scenario where nodes fail, along with all their associated edges. Failures could also be analyzed at the level of individual edges rather than entire nodes. The impact of single-edge failures on network substructures would vary based on the criteria used for edge selection, potentially leading to diverse outcomes in terms of network performance and functionality. By not focusing on individual edge removals, we concentrate on assessing the network’s robustness when confronted with the complete loss of connections from a critical node. This deliberate choice allows us to explore a severe form of network impairment, providing deeper insights into the resilience of decentralized learning under extreme conditions. We believe this approach offers a compelling analysis for initial robustness studies, which can later be refined, as suggested in Section 6, with more varied and nuanced failure dynamics.
@@
4.2.2 Time of disruption
In disaster scenarios, disruptions present themselves in unpredictable ways. They can happen as soon as the collaborative learning process starts, in the middle, or when it is approaching its end. The loss of nodes is expected to have a different impact depending on the current progress of the learning process: weak when it happens in the final phase, stronger early on. In the latter case, the key question is whether the decentralised learning process is able to recover and catch up over time. As discussed in Section 5.1, different learning configurations might exhibit different temporal dynamics, with one being slower than another, even if their accuracy converges to the same value. To account for these different learning speeds and enable fair comparisons among the various scenarios considered, we thus force disruptions to happen once the average accuracy of the system has reached a certain target threshold. For example, in Section 5.2, we consider disruptions happening when the average accuracy in the system has reached 0.7, 0.75, 0.8. This guarantees that disruptions affect similarly “skilled” systems in the learning process at the time of disruption.